40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 14 23:14:24 2018, aaron, Update, OMC, OMC scanning/aligning script 
    Reply  Tue Nov 20 20:33:11 2018, aaron, Update, OMC, OMC scanning/aligning script pathwaysketch.pdfIMG_0094.JPGIMG_0097.JPG
       Reply  Mon Nov 26 15:43:16 2018, aaron, Update, OMC, OMC scanning/aligning script c1omc.png
          Reply  Thu Dec 6 11:16:28 2018, aaron, Update, OMC, OMC channels 
             Reply  Mon Dec 10 12:29:05 2018, aaron, Update, OMC, OMC channels 52DE723A-02A4-4C62-879B-7B0070AE8A00.jpeg545E5512-D003-408B-9F00-55F985966A16.jpegDFF34976-CC49-4E4F-BFD1-A197E2072A32.jpeg
                Reply  Mon Dec 10 19:47:06 2018, aaron, Update, OMC, OMC channels 
                   Reply  Tue Dec 11 13:42:44 2018, Koji, Update, OMC, OMC channels 
                   Reply  Tue Dec 11 13:48:04 2018, aaron, Update, OMC, OMC channels 
                      Reply  Thu Dec 13 22:24:21 2018, aaron, Update, OMC, OMC channels 
          Reply  Mon Dec 10 12:11:28 2018, aaron, Update, OMC, Aligning the OMC B9DCF55F-1355-410C-8A29-EE45D43A56A4.jpeg
             Reply  Tue Dec 11 14:24:18 2018, aaron, Update, OMC, Aligning the OMC 
                Reply  Tue Dec 11 22:50:07 2018, aaron, Update, OMC, Aligning the OMC 6x
                   Reply  Thu Dec 13 22:36:42 2018, aaron, Update, OMC, Aligning the OMC 
                      Reply  Fri Dec 14 13:02:24 2018, aaron, Update, OMC, Aligning the OMC 
                      Reply  Fri Dec 14 13:06:12 2018, aaron, Update, OMC, Aligning the OMC 
                         Reply  Mon Dec 17 20:45:40 2018, aaron, Update, OMC, Aligning the OMC IMG_5922.JPGIMG_5921.JPG
                            Reply  Tue Dec 18 18:13:32 2018, aaron, Update, OMC, OMC L HV Piezo driver tests (again) IMG_4576_copy.jpg
                               Reply  Wed Dec 19 00:12:46 2018, gautam, Update, OMC, 40m OMC DCC node 
Message ID: 14365     Entry time: Tue Dec 18 18:13:32 2018     In reply to: 14363     Reply to this: 14366
Author: aaron 
Type: Update 
Category: OMC 
Subject: OMC L HV Piezo driver tests (again) 

I tested the OMC-L HV driver box again, and made the following observations:

  • Drove the HV diff pins (2,7) with a 5V triangle wave
    • Observed that with a ~0.4V offset on the drive, the HV output (measured directly with a 10x probe) has a 0-(almost)200V triangle wave (for 200V HV in), saturating near 200V and near 0V somewhat before reaching the full range of the triangle
    • The HV mon gives the same answer as measuring the HV output directly, and is reduced 100x compared to the HV output.
    • At 1Hz and above, the rolloff of the low pass still attenuates the drive a bit, and we don't reach the full range.
  • Drove the HV dither pins (1,6) with a 100mV to 10V triangle wave, around 15kHz
    • Even at 10V, the dithering is near the noise of the mon channel, so while I could see a slight peak changing on the FFT near the dither frequency, I couldn't directly observe this on a scope using the mon channel
    • However, measuring the HV directly I do see the dither applied on top of the HV signal. The amplitude of the dither is the same on the HV output as on the dither drive.

[gautam, aaron]

We searched for blips while nominally scanning the OMC length.

We sent a 0.1Hz, 10Vpp triangle wave to the OMC piezo drive diff channels, so the piezo length is seeing a slow triangle wave from 0-200V.

Then, we applied a ~15kHz dither to the OMC length. This dither is added directly onto the HV signal, so the amplitude of the dither at the OMC is the same as the amplitude of the dither into the HV driver.

We monitored the OMC REFL signal (where we saw no blips yesterday) and mixed this with the 15kHz dither signal to get an error signal. Gautam found a pomona box with a low pass filter, so we also low passsed to get rid of some unidentified high frequency noise we were seeing (possibly a ground loop at the function generator? it was present with the box off, but gone with the AC line unplugged). [So we made our own lock-in amplifier.] Photo attached.

We tested the transfer function of the LP, and finding it at 100kHz rather than the advertised 10kHz, we opened the box, removed a resistor to change the 3dB back to 10kHz, and confirmed this by measuring the TF.

We didn't see flashes of error signal in the mixed reflection either, so we suspect that either the PZT is not actuating on the OMC or the alignment is bad. Based on what appears to be the shimmering of far-misaligned fringes on the AS camera, Aaron's suspicion from aligning the cavity with the card, and the lack of flashes, we suspect the alignment. To avoid being stymied by a malfunctioning PZT, we can scan the laser frequency next time rather than the PZT length.

Attachment 1: IMG_4576_copy.jpg  1.935 MB  Uploaded Tue Dec 18 22:05:39 2018  | Hide | Hide all
IMG_4576_copy.jpg
ELOG V3.1.3-